Jim has sold a hand-carved wooden statue for $670, making a 35% profit. How much did it cost Jim to make the statue?
Answer by checkley75(3666)

1.35x=670
x=670/1.35
x=496.30 is the price before the profit.
proof
496.30+.35*496.30=670
496.30+173.70=670
670=670